# Exploring the Frontier of Vision-Language Models: A Survey of Current Methodologies and Future Directions

### v1

@@ 2.3 Multimodal Output with Multimodal Input
The Composable Diffusion (CoDi): CoDi Tang et al. (2023a) model adopts a multimodal approach using Latent Diffusion Models for text, image, video, and audio. Text processing involves a variational encoder (VAE) with BERT and GPT-2, image tasks use a latent diffusion model (LDM) with a VAE, and audio tasks utilize an LDM with a VAE encoder-decoder for mel-spectrogram representation. CoDi creates a shared multimodal space through cross-modal generation with Joint Multimodal Generation and cross-attention modules. Training involves individual diffusion models with aligned prompt encoders, and CoDi achieves any-to-any generation with a linear number of training objectives.
@@
CoDi-2 Tang et al. (2023b): CoDi-2 employs a multimodal encoder, ImageBind, with aligned encoders and a multilayer perceptron for modalities projection. It integrates diffusion models (DMs) into the multimodal latent language model (MLLM) for detailed, modality-interleaved generation. The fusion strategy involves projecting multimodal data into a feature sequence, processed by the MLLM, and utilizing DMs for improved generation quality. The alignment method leverages projections from aligned multimodal encoders to enable the MLLM to understand modality-interleaved input sequences, facilitating in-context learning and supporting multi-round interactive conversations.
@@
Google Gemini Google et al. (2023): Gemini models feature a transformative architecture with deep fusion capabilities, excelling in integrating text, image, audio, and video modalities. They surpass GPT-4 in 30 out of 32 benchmarks and are trained on Google’s TPU v4 and v5e accelerators for efficient scaling. The multimodal and multilingual training dataset prioritizes quality and safety, with models undergoing Reinforcement Learning from Human Feedback (RLHF). While specific details remain undisclosed, safety evaluations for bias and toxicity are a central part of Gemini’s development, involving collaboration with external experts.
@@
NExT-GPT Wu et al. (2023b): NExT-GPT features three stages: Multimodal Encoding, LLM Understanding and Reasoning, and Multimodal Generation. It uses models like ImageBind for encoding and Transformer-based layers for generation. In inference, modal encoders transform inputs, LLM decides on content, and diffusion decoders use signal tokens for synthesis. The system employs Multimodal Alignment Learning to align features and Modality-switching Instruction Tuning (MosIT) for improved LLM capabilities by aligning modal signal tokens with gold captions. The diverse MosIT dataset enhances the model’s ability to handle various user interactions effectively.
@@
VideoPoet Wu et al. (2023b): VideoPoet is a language model designed for high-quality video synthesis with matching audio. The model employs a decoder-only transformer architecture, processing multimodal inputs like images, videos, text, and audio. Utilizing a two-stage training protocol, VideoPoet showcases state-of-the-art capabilities in zero-shot video generation and excels in tasks such as text-to-video and video stylization. Notable features include a Large Language Model backbone, custom spatial super-resolution, and scalability with model size. Human evaluations highlight VideoPoet’s superiority in text fidelity, video quality, and motion interestingness. Responsible AI analysis underscores considerations for fairness, emphasizing the model’s capabilities in zero-shot editing, task chaining, and maintaining quality across multiple stages in video generation.
@@ 3 Future Directions
Tradeoff between pre-training and modular structure: A Lot of research is going on to increase the understanding, control, and faithfulness capacities of VLMs by introducing modularity in place of black box pretraining. Incorporating other modalities :Works are going on for incorporating more finer modalities like gaze/gestures inspired by Cheng et al. (2022) which is very important for the educational sector.
Fine-grained Evaluation of VLMs: Works are going on for more fine-grained evaluation of VLMs on parameters like bias, fairness etc. DALL-Eval Cho et al. (2023a) and VP-Eval Cho et al. (2023b) are a few works in this direction.
Causality and Counterfactual Capabilities in VLMs: A lot of work is done to understand the causal and counterfactual capabilities of LLM, which inspired researchers to explore the same in the realm of VLMs. Cm3 Aghajanyan et al. (2022) is one of the earliest works in this domain and there is a lot of buzz in this topic.
Continual Learning/Unlearning: A trend is there on efficiently learning continuously without training it from scratch in the VLM space. VQACL Zhang et al. (2023a) and Decouple before Interact Qian et al. (2023) are some of the earliest works in this domain. Inspired by the concept of knowledge unlearning observed in LLMs Si et al. (2023), researchers are also delving into similar approaches within the realm of VLMs.
@@
Multilingual grounding of VLMs: Following the recent surge in multilingual LLMs such as OpenHathi sarvam.ai (2023) and BharatGPT corovor.ai (2023), there is a growing momentum towards the development of multilingual vision-language models (VLMs).
More Domain specific VLMs: Various domain-specific VLMs, exemplified by projects like MedFlamingo Moor et al. (2023) and SkinGPT Zhou et al. (2023), have paved the way in their specialized fields. Further endeavors are in progress to craft VLMs specifically tailored for sectors like education and agriculture.

### v2

@@ 2.3 Multimodal Output with Multimodal Input
The Composable Diffusion (CoDi): CoDi [78] model adopts a multimodal approach using Latent Diffusion Models for text, image, video, and audio. Text processing involves a variational encoder (VAE) with BERT and GPT-2, image tasks use a latent diffusion model (LDM) with a VAE, and audio tasks utilize an LDM with a VAE encoder-decoder for mel-spectrogram representation. CoDi creates a shared multimodal space through cross-modal generation with Joint Multimodal Generation and cross-attention modules. Training involves individual diffusion models with aligned prompt encoders, and CoDi achieves any-to-any generation with a linear number of training objectives.
@@
CoDi-2 [79]: CoDi-2 employs a multimodal encoder, ImageBind, with aligned encoders and a multilayer perceptron for modalities projection. It integrates diffusion models (DMs) into the multimodal latent language model (MLLM) for detailed, modality-interleaved generation. The fusion strategy involves projecting multimodal data into a feature sequence, processed by the MLLM, and utilizing DMs for improved generation quality. The alignment method leverages projections from aligned multimodal encoders to enable the MLLM to understand modality-interleaved input sequences, facilitating in-context learning and supporting multi-round interactive conversations.
@@
Google Gemini [31]: Gemini models feature a transformative architecture with deep fusion capabilities, excelling in integrating text, image, audio, and video modalities. They surpass GPT-4 in 30 out of 32 benchmarks and are trained on Google’s TPU v4 and v5e accelerators for efficient scaling. The multimodal and multilingual training dataset prioritizes quality and safety, with models undergoing Reinforcement Learning from Human Feedback (RLHF). While specific details remain undisclosed, safety evaluations for bias and toxicity are a central part of Gemini’s development, involving collaboration with external experts.
@@
NExT-GPT [90]: NExT-GPT features three stages: Multimodal Encoding, LLM Understanding and Reasoning, and Multimodal Generation. It uses models like ImageBind for encoding and Transformer-based layers for generation. In inference, modal encoders transform inputs, LLM decides on content, and diffusion decoders use signal tokens for synthesis. The system employs Multimodal Alignment Learning to align features and Modality-switching Instruction Tuning (MosIT) for improved LLM capabilities by aligning modal signal tokens with gold captions. The diverse MosIT dataset enhances the model’s ability to handle various user interactions effectively.
@@
VideoPoet [90]: VideoPoet is a language model designed for high-quality video synthesis with matching audio. The model employs a decoder-only transformer architecture, processing multimodal inputs like images, videos, text, and audio. Utilizing a two-stage training protocol, VideoPoet showcases state-of-the-art capabilities in zero-shot video generation and excels in tasks such as text-to-video and video stylization. Notable features include a Large Language Model backbone, custom spatial super-resolution, and scalability with model size. Human evaluations highlight VideoPoet’s superiority in text fidelity, video quality, and motion interestingness. Responsible AI analysis underscores considerations for fairness, emphasizing the model’s capabilities in zero-shot editing, task chaining, and maintaining quality across multiple stages in video generation.
@@ 3 Future Directions
Tradeoff between pre-training and modular structure: A Lot of research is going on to increase the understanding, control, and faithfulness capacities of VLMs by introducing modularity in place of black box pretraining. Incorporating other modalities :Works are going on for incorporating more finer modalities like gaze/gestures inspired by [18] which is very important for the educational sector.
Fine-grained Evaluation of VLMs: Works are going on for more fine-grained evaluation of VLMs on parameters like bias, fairness etc. DALL-Eval [19] and VP-Eval [20] are a few works in this direction.
Causality and Counterfactual Capabilities in VLMs: A lot of work is done to understand the causal and counterfactual capabilities of LLM, which inspired researchers to explore the same in the realm of VLMs. Cm3 [3] is one of the earliest works in this domain and there is a lot of buzz in this topic.
Continual Learning/Unlearning: A trend is there on efficiently learning continuously without training it from scratch in the VLM space. VQACL [109] and Decouple before Interact [69] are some of the earliest works in this domain. Inspired by the concept of knowledge unlearning observed in LLMs [73], researchers are also delving into similar approaches within the realm of VLMs.
@@
Multilingual grounding of VLMs: Following the recent surge in multilingual LLMs such as OpenHathi [72] and BharatGPT [21], there is a growing momentum towards the development of multilingual vision-language models (VLMs). Palo[61] is the first notable work in this direction .
More Domain specific VLMs: Various domain-specific VLMs, exemplified by projects like MedFlamingo [63] and SkinGPT [117], have paved the way in their specialized fields. Further endeavors are in progress to craft VLMs specifically tailored for sectors like education and agriculture.
@@ 4 Conclusion
5 Acknowledgements
Akash Ghosh and Sriparna Saha express their heartfelt gratitude to the SERB (Science and Engineering Research Board ) POWER scheme(SPG/2021/003801) of the Department of Science and Engineering, Govt. of India, for providing the funding for carrying out this research